STATE-TO-FIELD VIBRATIONAL ENERGY TRANSFER FROM S1 PARA-DIFLUOROBENZENE IN COL-LISIONS WITH SEVERAL INERT GAS PARTNERS.

\underline{T. A. STONE}, D. L. CATLETT, AND C. S. PARMENTER, Indiana University Department of Chemistry, Bloomington, Indiana, 47405..

State-to-field vibrational energy transfer (VET) from several levels within the S_{1} (^{1}B_{2u}) manifold of para-Difluorobenzene (pDFB) vapor at 300 K in collisions with a series of gases is currently being examined. Vibrational relaxation by collisions with H_{2}, He^{1}, Ne, Ar^{1}, Kr, and Xe from five initial levels with vibrational energies ranging to 1634 cm^{−1} is probed. The resulting fluorescence spectra allow for the determination of absolute rate constants and cross sections for the collisional transfer of vibrational energy from the prepared single vibronic level into the surrounding S_{1} vibronic field of states. 1 D. L. Catlett and C. S. Parmenter, J. Phys. Chem., 95, 2864 (1991).